OpenStudy (anonymous):

What is the midpoint of the line segment with endpoints (5, -2) and (1, -6)?

OpenStudy (anonymous):

\[(\frac{ x1+x2 }{ 2 }, \frac{ y1+y2 }{ 2 })\] \[\left( \frac{ 5+1 }{ 2 }, \frac{ -6+(-2) }{ 2 } \right)\]

OpenStudy (anonymous):

is it 3 and -4

OpenStudy (anonymous):

Good job :o)